In a group there are M men and W women; m of the men and w of the women are college graduates. An individual is picked at random. Let A be the event the individual is a woman and B be the event he or she is a college graduate. Under what condition are the events A and B independent?
Answer #1

Total number of people = M+W

P(Acap B)= P(woman and college graduate) =rac{w}{M+W}

P(A) P(woman) =rac{W}{M+W}

P(B) P(college graduate) =rac{m+w}{M+W}

In order for A and B to be independent, we should have P(Acap B)=P(A)P(B)

rac{w}{M+W}=rac{W}{M+W} imes rac{m+W}{M+W}

w(M+W) = W(m+w)

wM = WM

w = W

which is the required condition.